OpenStudy (anonymous):

What does it mean/ what function do I write if the prob says: Given that f(x) varies inversely with the square of x?

OpenStudy (anonymous):

Not sure what the question is really stating but a function f(x) that is inversely proportional to x is written as f(x) = 1/x ; so function related to the inverse square would be f(x) = 1/x^{2} hope this is what you are looking for

OpenStudy (whpalmer4):

\[f(x) = k/x^2\] where k is a constant. If x doubles, f(x) is divided by 4.

OpenStudy (whpalmer4):

The graph will have this general shape:

OpenStudy (whpalmer4):

If k < 0, the "funnel" will point in the other direction.

OpenStudy (anonymous):

thank you!!!